A culture of bacteria grew from 320 to 780 in 5 hours. If this growth continues at this rate then how long will it take the population to triple
1
Expert's answer
2011-07-21T09:39:08-0400
Bacteria culture grows according the law:
A = A0 e^kt
Let's construct the system of equations
780 = 320 e 5k
780*3 = 780 ekt

5k = ln(780/320)
k = 0.178

3 = e 0.178 t
t = ln 3 / 0.178 = 6.165 hours